# Understanding the Performance Gap in Preference Learning: A Dichotomy of RLHF and DPO

### v1

@@ C.10 Proofs of Theorems 10 and 11
$\log\underset{\omega\sim\pi_{{\mathsf{ref}}}(\cdot|y)}{\mathbb{E}}\exp(\psi(y,\omega)_{1})=\left(2\textbf{r}_{\textup{dense}}-\textbf{r}_{\textup{sparse}}\right)^{\top}\psi(y)+C_{5}\;,$
where $C_{5}\in\mathbb{R}$ is an offset. And this is realizable as long as:
$\max_{\omega}\psi(\omega)_{1}-\min_{\omega}\psi(\omega)_{1}\geq\max_{y}(\textbf{r}_{\textup{dense}}-\textbf{r}_{\textup{sparse}})^{\top}\psi(y)-\min_{y}(\textbf{r}_{\textup{dense}}-\textbf{r}_{\textup{sparse}})^{\top}\psi(y)\;,$
which can be taken as a mild assumption on the features because $\psi$ and $\textbf{r}_{\textup{dense}},\textbf{r}_{\textup{sparse}}$ are all bounded. Now we are ready to prove the result.
@@
we thus have $\theta^{\star}_{r,0}=\textbf{r}_{\textup{sparse}}$ while $\theta^{\star}_{p,0}=2\textbf{r}_{\textup{dense}}$, due to the non-singularity of the Gram matrix.
@@
$\hat{\theta}_{r,\textup{rel}\ell_{1}}$ $=\underset{\theta_{0}+\tau_{1}\in\Theta_{B},\theta_{1}=e_{1}}{\operatorname{argmin}}-\frac{1}{N}\sum_{i=1}^{N}\log\sigma(r_{\theta}(y_{w}^{(i)}y_{w}^{(i)})-r_{\theta}(y_{l}^{(i)}y_{l}^{(i)}))+\gamma\|\theta_{0}\|_{1}\;,$
$\implies\hat{\theta}_{r,\textup{rel}\ell_{1},0}$ $=\underset{\theta_{0}+\tau_{1}\in\Theta_{B}}{\operatorname{argmin}}-\frac{1}{N}\sum_{i=1}^{N}\log\sigma(\beta(\theta_{0}+\tau_{1})^{\top}(\psi(y_{w}^{(i)})-\psi(y_{l}^{(i)}))+\gamma\|\theta_{0}+\tau_{1}-\tau_{1}\|_{1}\;,$
where $\tau_{1}:=e_{1}+\textbf{r}_{\textup{dense}}$. Then Lemma 5 implies there exists appropriate $\gamma,\delta$, such that w.p. $1-\delta$,
$\|\hat{\theta}_{r,\textup{rel}\ell_{1},0}-\theta^{\star}_{r,0}\|_{\Sigma_{\mathcal{D}}}^{2}=\mathcal{O}\left(\sqrt{\frac{k\log(2d)+k\log(1/\delta)}{n}}\right)\;.$
Now observe the empirical MLE objective of surrogate reward learning:
$\hat{\theta}_{p}$ $=\underset{\theta_{0}+\tau_{2}\in\Theta_{B},\theta_{1}=e_{1}}{\operatorname{argmin}}-\frac{1}{N}\sum_{i=1}^{N}\log\sigma(\hat{r}_{\theta}(y_{w}^{(i)}y_{w}^{(i)})-\hat{r}_{\theta}(y_{l}^{(i)}y_{l}^{(i)}))\;,$
$\implies\hat{\theta}_{p,0}$ $=\underset{\theta_{0}+\tau_{2}\in\Theta_{B}}{\operatorname{argmin}}-\frac{1}{N}\sum_{i=1}^{N}\log\sigma(\beta(\theta_{0}+\tau_{2})^{\top}(\psi(y_{w}^{(i)})-\psi(y_{l}^{(i)})))\;,$
where $\tau_{2}:=e_{1}-\textbf{r}_{\textup{dense}}+\textbf{r}_{\textup{sparse}}$. Then we can directly apply Lemma 1 and obtain:
$\underset{e_{1}+\textbf{r}_{\textup{dense}}+\textbf{r}_{\textup{sparse}}\in\Theta_{B}}{\sup}\|\hat{\theta}_{p,0}-\theta^{\star}_{p,0}\|_{\Sigma_{\mathcal{D}}}^{2}=\Omega\left(\frac{d}{n}\right)\;.$
C.11 Proof of Lemma 5
@@
We take $\gamma=\Theta\left(\sqrt{\frac{\log 2d+\log(1/\delta)}{n}}\right)$, where the specific value of $\gamma$ and $\delta$ are determined in Theorem 3.3 of Yao et al. [2025]. By the definition of the relative $\ell_{1}$-regularized estimator, we have:
@@
Note that $\theta^{\star},\tau\in\Theta_{B}$, thus $\|\theta^{\star}-\tau\|_{2}=\mathcal{O}(B)$. Then by Cauchy-Schwartz inequality and the fact that $\|\theta^{\star}-\tau\|_{0}=k$, we have $\|\theta^{\star}-\tau\|_{1}=\mathcal{O}(B\sqrt{k})$, and finally:
$\|\hat{{\theta}}_{\textup{rel}\ell_{1}}-{\theta}^{\star}\|_{{\Sigma_{\mathcal{D}}}}^{2}=\mathcal{O}\left(\sqrt{\frac{k\log(2d)+k\log(1/\delta)}{n}}\right)\;.$
C.12 Omitted Calculations in Sections 3 and B
Calculation of the performance gap. Since $\pi^{\star}(y)=\pi_{{\mathsf{ref}}}(y)\exp(r(y)/\beta)/Z$, we have that:
and for DPO we also have:
@@ C.12 Omitted Calculations in Sections 3 and B
term 1 $=\underset{y\sim\pi_{\theta^{\star}(r_{\phi})}}{\mathbb{E}}\nabla_{\phi}\log\pi_{\theta^{\star}(r_{\phi})}(y)[r^{\star}(y)-r_{\phi}(y)]$
@@
term 2 $=\underset{y\sim\pi_{\theta^{\star}(r_{\phi})}}{\mathbb{E}}\nabla r_{\phi}(y)$
@@
and we can assign the value of $\sigma^{\prime}(r_{\phi}(y)-r_{\phi}(y^{\prime}))$ to the sampling probability $\mu(y,y^{\prime})$. Thus we expect $\mu(y,y^{\prime})\propto\pi_{\theta^{\star}(r_{\phi})}/\sigma^{\prime}(r_{\phi}(y)-r_{\phi}(y^{\prime}))$. And under the context of surrogate reward learning, we have $\pi_{\theta^{\star}(r_{\phi})}=\pi_{\theta}$ and $r_{\phi}=\hat{r}_{\theta}$, and thus $\mu\propto\pi_{\theta^{\star}(r_{\phi})}/\sigma^{\prime}(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime}))$, which is exactly PILAF sampler (see Definition 1).
@@
$\mathcal{L}_{\textup{IPO}}^{\textup{online}}(\theta)$ $=\underset{(y,y^{\prime})\sim\mathsf{sg}\left(\rho_{\theta}\right)}{\mathbb{E}}p^{\star}(y\succ y^{\prime})\left[(r_{\theta}(y)-r_{\theta}(y^{\prime}))-\frac{1}{2}\right]^{2}+p^{\star}(y^{\prime}\succ y)\left[(r_{\theta}(y^{\prime})-r_{\theta}(y))-\frac{1}{2}\right]^{2}\;,$
@@
$\quad\nabla_{\theta}\mathcal{L}_{\textup{IPO}}^{\textup{online}}(\theta)$
$=2\underset{(y,y^{\prime})\sim\mathsf{sg}\left(\rho_{\theta}\right)}{\mathbb{E}}\left\{p^{\star}(y\succ y^{\prime})\left[(r_{\theta}(y)-r_{\theta}(y^{\prime}))-\frac{1}{2}\right]+p^{\star}(y^{\prime}\succ y)\left[(r_{\theta}(y)-r_{\theta}(y^{\prime}))+\frac{1}{2}\right]\right\}\nabla_{\theta}(r_{\theta}(y)-r_{\theta}(y^{\prime})$
$=2\underset{(y,y^{\prime})\sim\mathsf{sg}\left(\rho_{\theta}\right)}{\mathbb{E}}\left[(r_{\theta}(y)-r_{\theta}(y^{\prime}))-\frac{p^{\star}(y\succ y^{\prime})-p^{\star}(y^{\prime}\succ y)}{2}\right]\nabla_{\theta}(r_{\theta}(y)-r_{\theta}(y^{\prime})\;,$
@@
$\mathcal{L}_{\textup{IPO}}^{\textup{online}}(\theta)$ ${\stackrel{{\scriptstyle\scriptstyle\nabla}}{{=}}}\underset{(y,y^{\prime})\sim\mathsf{sg}\left(\rho_{\theta}\right)}{\mathbb{E}}\left[(r_{\theta}(y)-r_{\theta}(y^{\prime}))-\frac{p^{\star}(y\succ y^{\prime})-p^{\star}(y^{\prime}\succ y)}{2}\right]^{2}\;.$
C.13 Omitted Calculations in Section 4
Calculation of the performance gap. Motivated by Nika et al. [2024] and assuming $\mathcal{F}\subseteq\mathcal{F}_{\Pi}$, we can split the performance gap of two-stage RLHF again as:
@@ C.13 Omitted Calculations in Section 4
$=\underbrace{\underset{y\sim\pi^{\star}}{\mathbb{E}}\left[r^{\star}(y)-\hat{r}_{\theta}(y)\right]+\underset{y\sim\pi_{\theta}}{\mathbb{E}}\left[\hat{r}_{\theta}(y)-r^{\star}(y)\right]}_{\textup{ implicit reward representation gap}}\;,$
$\pi^{\star}(y_{t}|y_{0\ldots t-1})=\pi_{{\mathsf{ref}}}(y_{t}|y_{0\ldots t-1})\exp\left(\frac{q(y_{t}|y_{0\ldots t-1})-q(y_{t-1}|y_{0\ldots t-2})}{\beta}\right)\;,$
where the $q$ function is determined in a recursive way:
$q(y_{t}|y_{0\ldots t-1})=\begin{cases}\beta\log\sum_{s}\pi_{{\mathsf{ref}}}(s|y_{0\ldots t})\exp(q(s|y_{0\ldots t})/\beta)&y_{t}\text{ is not the terminal token;}\\ r^{\star}(y_{0\ldots t})&y_{t}\text{ is the terminal token.}\end{cases}$
To prove this, we define a $q^{\prime}$ function as:
$q^{\prime}(y_{0})=Z+\beta\log\frac{\pi^{\star}(y_{0})}{\pi_{{\mathsf{ref}}}(y_{0})}\;,\;q^{\prime}(y_{t}|y_{0\ldots t-1})=q^{\prime}(y_{t-1}|y_{0\ldots t-2})+\beta\log\frac{\pi^{\star}(y_{t}|y_{0\ldots t-1})}{\pi_{{\mathsf{ref}}}(y_{t}|y_{0\ldots t-1})}\;,$
where $Z:=\log\sum_{y}\pi_{{\mathsf{ref}}}(y)\exp(r^{\star}(y)/\beta)$. And then for a $y$ with $y_{N}$ as the terminal token, we have:
$=q^{\prime}(y_{0})-Z+\sum_{t=1}^{N}q^{\prime}(y_{t}|y_{0\ldots t-1})-q^{\prime}(y_{t-1}|y_{0\ldots t-2})$
$=-Z+q^{\prime}(y_{N}|y_{0\ldots N-1})\;.$
$q^{\prime}(y_{N}|y_{0\ldots N-1})=r^{\star}(y)\;.$ (7)
Then by definition:
we have:
$\sum_{y}\pi_{{\mathsf{ref}}}(y)\exp\left(\frac{q^{\prime}(y_{t}|y_{0\ldots t-1})-q^{\prime}(y_{t-1}|y_{0\ldots t-2})}{\beta}\right)=1\;,$
which yields:
$q^{\prime}(y_{t-1}|y_{0\ldots t-2})=\beta\log\sum_{s}\pi_{{\mathsf{ref}}}(s|y_{0\ldots t-1})\exp(q(s|y_{0\ldots t-1})/\beta)\;.$ (8)
Combining Equations 7 and 8, we complete the calculations.
Appendix D Implementation Details
Codebases. Our codebase is mainly based on MODPO [Zhou et al., 2024] (https://github.com/ZHZisZZ/modpo), Online-RLHF [Dong et al., 2024, Xiong et al., 2024] (https://github.com/RLHFlow/Online-RLHF), Samplers-in-Online-DPO [Shi et al., 2025] (https://github.com/srzer/Samplers-in-Online-DPO). We are committed to releasing the codes.
Datasets. We adopt one common training datasets, PKU-SafeRLHF [Ji et al., 2023] (https://huggingface.co/datasets/PKU-Alignment/PKU-SafeRLHF). SFT: We train our initial model on $5$k samples of PKU-SafeRLHF-QA (https://huggingface.co/datasets/PKU-Alignment/PKU-SafeRLHF-QA). Online training: We use $10$k samples of PKU-SafeRLHF-Prompt (https://huggingface.co/datasets/PKU-Alignment/PKU-SafeRLHF-prompt) for training, and $2$k samples for evaluation. Offline training: We adopt two preference datasets, PKU-SafeRLHF-safer and PKU-SafeRLHF-better, each composed of $9$k training samples and $2$k evaluation samples, following the practice of Zhou et al. [2024].

### v2

@@ B.10.3 Proof of Theorem 13
$\log\underset{\omega\sim\pi_{{\mathsf{ref}}}(\cdot|y)}{\mathbb{E}}\exp(\psi(y,\omega)_{1})=\textbf{r}_{\textup{dense}}^{\top}\psi(y)+C_{5}\;,$
for $\forall y$, where $C_{5}\in\mathbb{R}$ is an offset.
@@
we thus have $\theta^{\star}_{r,0}=\textbf{r}_{\textup{sparse}}$ and $\theta^{\star}_{p,0}=\textbf{r}_{\textup{sparse}}+\textbf{r}_{\textup{dense}}$, due to the non-singularity of the Gram matrix.
@@
$\hat{\theta}_{r,\textup{rel}\ell_{1}}$ $\in\underset{\theta_{0}+\tau_{1}\in\Theta_{B},\theta_{1}=e_{1}}{\operatorname{argmin}}-\frac{1}{n}\sum_{i=1}^{n}\log\sigma(r_{\theta}(\bar{y}_{w}^{(i)}\bar{y}_{w}^{(i)})-r_{\theta}(\bar{y}_{l}^{(i)}\bar{y}_{l}^{(i)}))+\gamma\|\theta_{0}\|_{1}\;,$
$\implies\hat{\theta}_{r,\textup{rel}\ell_{1},0}$ $\in\underset{\theta_{0}+\tau_{1}\in\Theta_{B}}{\operatorname{argmin}}-\frac{1}{n}\sum_{i=1}^{n}\log\sigma(\beta(\theta_{0}+\tau_{1})^{\top}(\psi(\bar{y}_{w}^{(i)})-\psi(\bar{y}_{l}^{(i)}))+\gamma\|\theta_{0}+\tau_{1}-\tau_{1}\|_{1}\;,$
where $\tau_{1}:=e_{1}+\textbf{r}_{\textup{dense}}$. Then Lemma 4 implies there exists appropriate $\gamma$, such that w.p. $1-\delta$,
$\frac{1}{n}\sum_{i=1}^{n}\left[(\hat{\theta}_{r,\textup{rel}\ell_{1},0}-\textbf{r}_{\textup{sparse}})^{\top}(\psi(\bar{y}_{w}^{(i)})-\psi(\bar{y}_{l}^{(i)}))\right]^{2}=\mathcal{O}\left(\sqrt{\frac{k\log(d)+k\log(1/\delta)}{n}}\right)\;,\;$
and thus w.p. $1-\delta$,
$\quad\frac{1}{n}\sum_{i=1}^{n}\left[(r^{\star}(y_{w}^{(i)})-r^{\star}(y_{l}^{(i)}))-(r_{\hat{\theta}_{r,\textup{rel}\ell_{1}}}(y_{w}^{(i)})-r_{\hat{\theta}_{r,\textup{rel}\ell_{1}}}(y_{l}^{(i)}))\right]^{2}$
$=\frac{\beta^{2}}{n}\sum_{i=1}^{n}\left[(\textbf{r}_{\textup{sparse}}+\textbf{r}_{\textup{dense}}+e_{1})^{\top}(\psi(\bar{y}_{w}^{(i)})-\psi(\bar{y}_{l}^{(i)}))-(\theta_{r,\textup{rel}\ell_{1},0}+\textbf{r}_{\textup{dense}}+e_{1})^{\top}(\psi(\bar{y}_{w}^{(i)})-\psi(\bar{y}_{l}^{(i)}))\right]^{2}$
$=\frac{\beta^{2}}{n}\sum_{i=1}^{n}\left[(\textbf{r}_{\textup{sparse}}-\theta_{r,\textup{rel}\ell_{1},0})^{\top}(\psi(\bar{y}_{w}^{(i)})-\psi(\bar{y}_{l}^{(i)}))\right]^{2}$
$=\mathcal{O}\left(\sqrt{\frac{k\log(d)+k\log(1/\delta)}{n}}\right)\;.$
Note that
$\log\sigma(\hat{r}_{\theta}(y_{w}^{(i)}y_{w}^{(i)})-\hat{r}_{\theta}(y_{l}^{(i)}y_{l}^{(i)}))=\log\sigma(\beta(\theta_{p,0}+e_{1})^{\top}(\psi(y_{w}^{(i)})-\psi(y_{l}^{(i)})))\;,$
then Lemma 1 implies that $\exists N=\mathcal{O}(d)$, for $\forall n\geq N$ and any estimator $\hat{\theta}_{p}$, we have
$\underset{e_{1}+\textbf{r}_{\textup{dense}}+\textbf{r}_{\textup{sparse}}\in\Theta_{B}}{\sup}\frac{1}{n}\sum_{i=1}^{n}\left[(\hat{\theta}_{p,0}+e_{1}-\textbf{r}_{\textup{sparse}}-\textbf{r}_{\textup{dense}}-e_{1})^{\top}(\psi(\bar{y}_{w}^{(i)})-\psi(\bar{y}_{l}^{(i)}))\right]^{2}=\Omega\left(\frac{d}{n}\right)\;.$
Now observe the data-induced semi-norm of surrogate reward learning:
$\quad\frac{1}{n}\sum_{i=1}^{n}\left[(r^{\star}(y_{w}^{(i)})-r^{\star}(y_{l}^{(i)}))-(r_{\hat{\theta}_{p}}(y_{w}^{(i)})-r_{\hat{\theta}_{p}}(y_{l}^{(i)}))\right]^{2}$
$=\frac{\beta^{2}}{n}\sum_{i=1}^{n}\left[(\textbf{r}_{\textup{sparse}}+\textbf{r}_{\textup{dense}}+e_{1})^{\top}(\psi(\bar{y}_{w}^{(i)})-\psi(\bar{y}_{l}^{(i)}))-(\hat{\theta}_{p}+e_{1})^{\top}(\psi(\bar{y}_{w}^{(i)})-\psi(\bar{y}_{l}^{(i)}))\right]^{2}$
$=\frac{\beta^{2}}{n}\sum_{i=1}^{n}\left[(\hat{\theta}_{p,0}+e_{1}-\textbf{r}_{\textup{sparse}}-\textbf{r}_{\textup{dense}}-e_{1})^{\top}(\psi(\bar{y}_{w}^{(i)})-\psi(\bar{y}_{l}^{(i)}))\right]^{2}\;.$
And thus there exists an environment for DTSP, s.t.
$\frac{1}{n}\sum_{i=1}^{n}\left[(r^{\star}(y_{w}^{(i)})-r^{\star}(y_{l}^{(i)}))-(r_{\hat{\theta}_{p}}(y_{w}^{(i)})-r_{\hat{\theta}_{p}}(y_{l}^{(i)}))\right]^{2}=\Omega\left(\frac{d}{n}\right)\;.$
B.10.4 Proof of Lemma 4
@@
We take $\gamma=\Theta\left(\sqrt{\frac{\log(d)+\log(1/\delta)}{n}}\right)$, where the specific value of $\gamma$ is determined in Theorem 3.3 of Yao et al. [2025]. By the definition of the relative $\ell_{1}$-regularized estimator, we have:
@@
Note that $\theta^{\star},\tau\in\Theta_{B}$, thus $\|\theta^{\star}-\tau\|_{2}=\mathcal{O}(1)$. Then by Cauchy-Schwartz inequality and the fact that $\|\theta^{\star}-\tau\|_{0}=k$, we have $\|\theta^{\star}-\tau\|_{1}=\mathcal{O}(\sqrt{k})$, and finally:
$\|\hat{{\theta}}_{\textup{rel}\ell_{1}}-{\theta}^{\star}\|_{{\Sigma_{\mathcal{D}}}}^{2}=\mathcal{O}\left(\sqrt{\frac{k\log(d)+k\log(1/\delta)}{n}}\right)\;.$
B.11 Omitted Calculations
$\pi^{\star}(y_{t}|y_{0\ldots t-1})$ $=\pi_{{\mathsf{ref}}}(y_{t}|y_{0\ldots t-1})\exp\left(\frac{q^{\star}(y_{t}|y_{0\ldots t-1})-q^{\star}(y_{t-1}|y_{0\ldots t-2})}{\beta}\right)\;,$
$\pi^{\star}(y_{0})$ $=\pi_{{\mathsf{ref}}}(y_{0})\exp\left(\frac{q^{\star}(y_{0})-\beta\log Z}{\beta}\right)\;,$
where $Z:=\sum_{y}\pi_{{\mathsf{ref}}}(y)\exp(r^{\star}(y)/\beta)$, and the $q^{\star}$ function is determined in a recursive way:
[…]
To prove this, we define a $q^{\prime}$ function as:
$q^{\prime}(y_{0})=\beta\log Z+\beta\log\frac{\pi^{\star}(y_{0})}{\pi_{{\mathsf{ref}}}(y_{0})}\;,\;q^{\prime}(y_{t}|y_{0\ldots t-1})=q^{\prime}(y_{t-1}|y_{0\ldots t-2})+\beta\log\frac{\pi^{\star}(y_{t}|y_{0\ldots t-1})}{\pi_{{\mathsf{ref}}}(y_{t}|y_{0\ldots t-1})}\;.$
For the initial token, by definition we have:
$\pi^{\star}(y_{0})$ $=\pi_{{\mathsf{ref}}}(y_{0})\exp\left(\frac{q^{\prime}(y_{0})-\beta\log Z}{\beta}\right)\;.$ (9)
And then for a $y$ with $y_{N}$ as the terminal token, we have:
$=q^{\prime}(y_{0})-\beta\log Z+\sum_{t=1}^{N}q^{\prime}(y_{t}|y_{0\ldots t-1})-q^{\prime}(y_{t-1}|y_{0\ldots t-2})$
$=-\beta\log Z+q^{\prime}(y_{N}|y_{0\ldots N-1})\;.$
$\beta\log\frac{\pi^{\star}(y)}{\pi_{{\mathsf{ref}}}(y)}=-\beta\log Z+r^{\star}(y)\;,$
thus
$q^{\prime}(y_{N}|y_{0\ldots N-1})=r^{\star}(y)\;.$ (10)
Then by definition:
we have:
$\pi_{{\mathsf{ref}}}(y_{t}|y_{0\ldots t-1})\exp\left(\frac{q^{\prime}(y_{t}|y_{0\ldots t-1})-q^{\prime}(y_{t-1}|y_{0\ldots t-2})}{\beta}\right)=\pi^{\star}(y_{t}|y_{0\ldots t-1})\;,$ (11)
and thus
$\sum_{s}\pi_{{\mathsf{ref}}}(s|y_{0\ldots t-1})\exp\left(\frac{q^{\prime}(y_{t}|y_{0\ldots t-1})-q^{\prime}(y_{t-1}|y_{0\ldots t-2})}{\beta}\right)=1\;,$
which yields:
$q^{\prime}(y_{t-1}|y_{0\ldots t-2})=\beta\log\sum_{s}\pi_{{\mathsf{ref}}}(s|y_{0\ldots t-1})\exp(q^{\prime}(s|y_{0\ldots t-1})/\beta)\;.$ (12)
Combining Equations 9, 11, 10 and 12, we show that $q^{\star}$ exists and is equivalent to $q^{\prime}$.
@@ B.11 Omitted Calculations
term 1
$=\underset{y\sim\pi_{\theta^{\star}(r_{\phi})}}{\mathbb{E}}\nabla_{\phi}\log\pi_{\theta^{\star}(r_{\phi})}(y)[r^{\star}(y)-r_{\phi}(y)]$
@@
term 2
$=\underset{y\sim\pi_{\theta^{\star}(r_{\phi})}}{\mathbb{E}}\nabla r_{\phi}(y)$
@@
and we can assign the value of $\sigma^{\prime}(r_{\phi}(y)-r_{\phi}(y^{\prime}))$ to the sampling probability $\mu(y,y^{\prime})$. Thus we expect $\mu(y,y^{\prime})\;\propto\;\pi_{\theta^{\star}(r_{\phi})}/\sigma^{\prime}(r_{\phi}(y)-r_{\phi}(y^{\prime}))$. And under the context of DPO, we have $\pi_{\theta^{\star}(r_{\phi})}=\pi_{\theta}$ and $r_{\phi}=\hat{r}_{\theta}$, and thus $\mu\;\propto\;\pi_{\theta^{\star}(r_{\phi})}/\sigma^{\prime}(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime}))$, which is exactly PILAF sampler.
@@
$\mathcal{L}_{\textup{IPO}}^{\textup{online}}(\pi_{\theta})$ $=\underset{(y,y^{\prime})\sim\mathsf{sg}\left(\rho_{\theta}\right)}{\mathbb{E}}p^{\star}(y\succ y^{\prime})\left[(r_{\theta}(y)-r_{\theta}(y^{\prime}))-\frac{1}{2}\right]^{2}+p^{\star}(y^{\prime}\succ y)\left[(r_{\theta}(y^{\prime})-r_{\theta}(y))-\frac{1}{2}\right]^{2}\;,$
@@
$\quad\nabla_{\theta}\mathcal{L}_{\textup{IPO}}^{\textup{online}}(\pi_{\theta})$
$=2\underset{(y,y^{\prime})\sim\mathsf{sg}\left(\rho_{\theta}\right)}{\mathbb{E}}\left\{p^{\star}(y\succ y^{\prime})\left[(r_{\theta}(y)-r_{\theta}(y^{\prime}))-\frac{1}{2}\right]+p^{\star}(y^{\prime}\succ y)\left[(r_{\theta}(y)-r_{\theta}(y^{\prime}))+\frac{1}{2}\right]\right\}\nabla_{\theta}(r_{\theta}(y)-r_{\theta}(y^{\prime}))$
$=2\underset{(y,y^{\prime})\sim\mathsf{sg}\left(\rho_{\theta}\right)}{\mathbb{E}}\left[(r_{\theta}(y)-r_{\theta}(y^{\prime}))-\frac{p^{\star}(y\succ y^{\prime})-p^{\star}(y^{\prime}\succ y)}{2}\right]\nabla_{\theta}(r_{\theta}(y)-r_{\theta}(y^{\prime}))\;,$
@@
$\mathcal{L}_{\textup{IPO}}^{\textup{online}}(\pi_{\theta})$ ${\stackrel{{\scriptstyle\scriptstyle\nabla}}{{=}}}\underset{(y,y^{\prime})\sim\mathsf{sg}\left(\rho_{\theta}\right)}{\mathbb{E}}\left[(r_{\theta}(y)-r_{\theta}(y^{\prime}))-\frac{p^{\star}(y\succ y^{\prime})-p^{\star}(y^{\prime}\succ y)}{2}\right]^{2}\;.$
Below we provide two additional results for reference.
Exact Optimization: calculation of the policy performance gap. Since the optimal policy satisfies $\pi^{\star}(y)=\pi_{{\mathsf{ref}}}(y)\exp(r(y)/\beta)/Z$, we have that:
and for DPO we also have:
Approximate Optimization: calculation of the reward representation gap. Motivated by Nika et al. [2024] and assuming $\mathcal{F}\subseteq\mathcal{F}_{\Pi}$, we can split the performance gap of two-stage RLHF again as:
@@
$=\underbrace{\underset{y\sim\pi^{\star}}{\mathbb{E}}\left[r^{\star}(y)-\hat{r}_{\theta}(y)\right]+\underset{y\sim\pi_{\theta}}{\mathbb{E}}\left[\hat{r}_{\theta}(y)-r^{\star}(y)\right]}_{\textup{ implicit reward representation gap}}\;.$
Appendix C Implementation Details
Codebases. Our codebase is mainly based on MODPO [Zhou et al., 2024] (https://github.com/ZHZisZZ/modpo), Online-RLHF [Dong et al., 2024, Xiong et al., 2024] (https://github.com/RLHFlow/Online-RLHF), Samplers-in-Online-DPO [Shi et al., 2025] (https://github.com/srzer/Samplers-in-Online-DPO). Our code is released at https://github.com/srzer/Gap-in-Preference-Learning.
Datasets. We adopt one common training dataset, PKU-SafeRLHF [Ji et al., 2023] (https://huggingface.co/datasets/PKU-Alignment/PKU-SafeRLHF). SFT: We train our initial model on 55k samples of PKU-SafeRLHF-QA (https://huggingface.co/datasets/PKU-Alignment/PKU-SafeRLHF-QA). Online training: We use 1010k samples of PKU-SafeRLHF-Prompt (https://huggingface.co/datasets/PKU-Alignment/PKU-SafeRLHF-prompt) for training, and 22k samples for evaluation. Offline training: We adopt two preference datasets, PKU-SafeRLHF-safer and PKU-SafeRLHF-better, each composed of 99k training samples and 22k evaluation samples, following the practice of Zhou et al. [2024].